OpenStudy (anonymous):

what is the y intercept of the equation 2x+y=7?

Directrix (directrix):

7

Directrix (directrix):

y = mx+b where m is the slope and b is the y-intercept 2x + y = 7 y = 7 - 2x y = -2x + 7 So, the slope is -2 and the y-intercept is 7. The y-intercept has coordinates (0,7).

OpenStudy (anonymous):

For y intercept, take x as 0 2(0)+y=7 y=7 y-intercept is 7.